OpenStudy (anonymous):

expressions that are equivalent to −3.5(2 − 3n) − 2.5n?

OpenStudy (anonymous):

−3.5(2 − 3n) − 2.5n = -7 +10.5n -.2.5n = -7 +8n

OpenStudy (anonymous):

thank you